OpenStudy (anonymous):

All I do is help people on openstudy and no one ever helps me with my work pissed off :*

OpenStudy (imstuck):

Do you need help with something?

OpenStudy (imstuck):

I experienced that same thing...but then I had to stop and think for a minute...I'm much more capable of helping others with the types of questions they have, while the types of questions I have are much too difficult for other to help me with. Don't be offended...maybe no one on OS at that time is knowledgeable enough to help.

OpenStudy (anonymous):

@Frazier2k you should post this in open study feedback

OpenStudy (anonymous):

if u need help u can ask me. but I cant promise I will be able too. im good at math but I still got a lot to learn. and im sorry ppl r not willing to help u. @Frazier2k

OpenStudy (anonymous):

you should tag ppl on your questions i can help

OpenStudy (anonymous):

ok i have one question

OpenStudy (anonymous):

The function f(x) = 245(4)x represents the growth of a fruit fly population every year in the orange grove. Stacey wants to manipulate the formula to an equivalent form that calculates every month, not every year. Which function is correct for Stacey's purposes? f(x) = 45(4)x f(x) = 45(412)the x over 12 power f(x) = 2752(4)x f(x) = 245(4 to the one twelfth power)12x

OpenStudy (anonymous):

I got the second answer choice

OpenStudy (anonymous):

is it the one u posted like 30 min ago?? I just saw tht one im sorry but I am unable to help since im just learning tht right now. I wish I could tho. sorry

OpenStudy (anonymous):

thats ok i think my answer is correct i just want to be reassured

OpenStudy (anonymous):

In the formula x is in years. If x should be in months, then replace x in the formula with x/12 and now x will be in months.

OpenStudy (anonymous):

Oh yeah @IMStuck i appreciate that bro you opened my eyes to another point of view

OpenStudy (anonymous):

thank u for the medal

OpenStudy (anonymous):

No Ty for the help

OpenStudy (anonymous):

yw